7) Tickets for a play at the community theater cost $20 for an adult and $8 for a child. If 220 tickets were sold and the total receipts were $3200, how many of each type of ticket were sold?
Can someone please help?
v/r
Jess

Answer by stanbon(75887)

Tickets for a play at the community theater cost $20 for an adult and $8 for a child. If 220 tickets were sold and the total receipts were $3200, how many of each type of ticket were sold?
----------
Using two variables
Quantity Equation: a + c = 220
Value EQUATION: 20a + 8c = 3200
---------------------
Modify to use elimination:
20a + 20c = 4400
20a + 8c = 3200
-------------------
Subtract to solve for "c"
12c = 1200
c = 100 (# of child tickets)
--------
Substitute to solve for "a":
a+c = a+100 = 220
a = 120 (# of adult tickets)
=============
Cheers,
Stan H.